# A decade of the fast-varying ionospheric and magnetospheric magnetic fields

### v1

@@ 5.3 Fourier spectrum of external fields
In this section we explore the frequency content of the derived time series of the external coefficients describing ionospheric and magnetospheric sources. Our 10-year time series of the SH coefficients has a theoretical bandwidth of over four decades, i.e. $f_{\mathrm{Nyq}}/\Delta f>10^{4}$, where $f_{\mathrm{Nyq}}$ is the Nyquist frequency and $\Delta f$ is the frequency resolution. As such, it covers the entire period range from sub-diurnal to inter-annual periods.
@@
The first zonal harmonic of the magnetospheric magnetic field, largely attributed to the ring current, is the most energetic contribution to the magnetic field of the external origin at periods longer than one day. The power of the spectrum for the first zonal harmonic quickly rolls off at periods shorter than one day, diurnal frequency bands, while exhibits strong amplitude at longer periods (Fig. 17, upper panel). At longest periods, a prominent spectral peak can be observed between $2.0$ and $2.5$ years. This biennial oscillation has been previously observed in the solar magnetic field, and has been attributed to a harmonic of the solar cycle (Knaack & Stenflo, 2005). This notable feature can be similarly observed in the ionospheric signal (Fig. 17 lower panel) and in other degree-one modes (Fig. LABEL:fig:spec-q11-mi). The second, rather sharp peak in the magnetospheric field spectrum is at the semi-annual period (2 cycles per year). Since our estimation is conducted in the Earth-fixed frame, the semi-annual peak is likely associated with the axial tilt of the Earth and the associated seasonal variability in the Earth’s frame (Maus & Lühr, 2005; Olsen et al., 2005; Balasis & Egbert, 2006). A secondary, wider peak can be observed at the annual period (1 cycle per year). This can be a signature of the tail current in the magnetosphere (Maus & Lühr, 2005). As the $q_{10}$ component experiences semi-annual changes in the solar radiation, it is further modulated by the change in the Interplanetary Magnetic Field. Both annual and semi-annual peaks are observed not only in the first zonal harmonic component, but also in other dipole components of the magnetospheric field (Fig. LABEL:fig:spec-q11-mi upper panel). This can result from the transformation of the quasi-static magnetospheric field in GSM/SM frames to the Earth-fixed frame. In contrast, these peaks are weaker or not visible in the ionospheric spectrum (Figs. 17 and LABEL:fig:spec-q11-mi, lower panels), suggesting that ionospheric dynamics at $n=1$ is not as strongly modulated by the orbit-related periodicities.
@@
Figure 17: Amplitude spectrum of the first zonal harmonic component $P_{1}^{0}$ of the external magnetic fields. The magnetospheric field ($q_{10}^{\mathrm{mag}}$) and ionospheric field ($q_{10}^{\mathrm{ion}}$) spectra are shown in the upper and the lower panels, respectively. Major spectral peaks observed in the spectra are marked with blue triangles, with their peak periods annotated. Selected theoretical periods of the system are plotted as gray vertical lines for reference.
@@
In contrast to variations at long periods, the variations in diurnal period band are dominated by the ionospheric sources. One of the major features in the ionospheric signal in this period band is the diurnal mode and its harmonics. These are clearly visible as discrete spectral peaks at one, two, three and four cycles per day (cpd) (Fig. 18, lower panel). For the $P_{2}^{1}$ spatial mode, the fundamental time harmonic with a period of one day is the strongest, with an amplitude two orders of magnitudes above the average spectrum level. This mode, which has an azimuthal wavenumber $m=1$, and has a 1-cpd periodicity observed in the Earth-fixed geographic frame, represents a stationary $n=2$ mode in a solar-fixed frame (local-time term). This mode is predominantly influenced by the day-side structure of the Sq current system (Malin, 1973; Winch, 1981; Schmucker, 1999). The same holds also for the SH coefficient $s_{32}^{\mathrm{ion}}$ and $s_{43}^{\mathrm{ion}}$ with the azimuthal wavenumbers of $m=2$ and $m=3$, respectively. They exhibit strongest spectral peaks at 2 cpd (Fig. LABEL:fig:spec-s32-mi lower panel) and 3 cpd (Fig. LABEL:fig:spec-s43-mi lower panel), respectively. These correspond to the $n=3$ and $n=4$ modes that are stationary in a solar-fixed frame. The diurnal periods are also visible in the spectra of magnetospheric coefficients (Figs. 18, LABEL:fig:spec-s32-mi and LABEL:fig:spec-s43-mi upper panels), although they are much weaker compared to the ionosphere.
Figure 18: Amplitude spectrum of the magnetospheric $q_{21}^{\mathrm{mag}}$ component (upper panel) and ionospheric $q_{21}^{\mathrm{ion}}$ component (lower panel). Spectral peaks observed in the spectra are marked with blue triangles, with their peak period annotated. Selected theoretical periods of the system, including the diurnal period and its harmonics, and the periods for lunar daily variations $L_{p}$ with $p=1,2,3,4$, are plotted as gray vertical lines for reference.
@@ 6 Discussion and Conclusions
Compared to the previous geomagnetic field models, our new modelling approach with the ionosphere parametrisation has several distinct features. First, the magnetic fields with ionospheric and magnetospheric origins are separated. This sets our model apart from the modelling approaches where the ionospheric field is neglected (Lesur et al., 2010; Finlay et al., 2020) or the magnetospheric field is removed by baseline removal and ring current corrections (Yamazaki & Maute, 2017; Guzavina et al., 2019). Second, no temporal behaviour is prescribed on either of the external fields. This is in contrast to many other models. For instance, the temporal behaviour for magnetospheric fields beyond $q_{10}$ is prescribed to be standing modes in SM or GSM coordinates in the CHAOS model (Finlay et al., 2020), or the ionospheric field is assumed to consist of daily and seasonal time-harmonics in the Comprehensive Model (Sabaka et al., 2018). Furthermore, the secondary induced field in the solid Earth is co-estimated alongside the external fields, rather than modelled using an existing conductivity model, as is the case in Chulliat et al. (2016).
@@
When the geomagnetic field model fails to capture the true ionospheric signal properly, either due to negligence or limited expressivity of the parametrisation of the ionospheric field, the remnant ionospheric signal might be compensated by biasing the induced field and the magnetospheric field estimates. To test whether the induced and the magnetospheric fields are indeed biased by the negligence of the ionospheric field, we need a physics-based metric to examine how the model with an ionosphere compares with a model without the ionosphere. For this purpose, we estimated the transfer functions relating the induced and the external parts of the geomagnetic potential. In the frequency domain, assuming a radially symmetric electrical conductivity of the Earth, the Gauss coefficients of the induced field are related to those of the external field as
@@
where $Q_{n}$ is the so-called $Q$-response (Schmucker, 1985), and is a functional of the subsurface conductivity $\sigma$. For a general 3-D subsurface conductivity, the relation above is not complete since a single external mode would induce many internal modes, all of which are coupled through a so-called Q-matrix (Olsen, 1999a). As the subsurface conductivity can be approximated to the first order as a radial conductivity profile, the diagonal elements of this matrix dominate, and the relation between the induced and external coefficients can to first order be approximated by the expression above. Therefore, the equation above is still a valid test since any realistic 3-D conductivity effects will be dwarfed by much stronger source effects.
@@
$\varepsilon_{nm}=\left\{\begin{aligned} &\frac{1}{2}\left(q_{n|m|}-is_{n|m|}\right),\quad m\geq 0,\\ &\frac{1}{2}\left(q_{n|m|}+is_{n|m|}\right),\quad m<0,\end{aligned}\right.\qquad\iota_{nm}=\left\{\begin{aligned} &\frac{1}{2}\left(g_{n|m|}-ih_{n|m|}\right),\quad m\geq 0,\\ &\frac{1}{2}\left(g_{n|m|}+ih_{n|m|}\right),\quad m<0.\end{aligned}\right.$ (19)
@@
The real part of the $C_{n}$-response is a depth of the maximum induced current, thus approximating the sounding depth at a given frequency (Weidelt, 1972). We estimated the transfer functions by performing robust regression on the spectral transformed induced and external Gauss coefficients (19). The resulting $C_{n}$-responses from the field estimates of the model with and without the ionosphere, along with the squared coherence, are visualised in Fig. 19 at $9$ log-spaced periods between $1.5$ days and $10$ days, as well as at the diurnal ($24$ hours), semidiurnal ($12$ hours) and terdiurnal ($8$ hours) periods. It is also possible to estimate these transfer functions at longer periods, but the effect of the ionosphere is less pronounced.
@@
The TFs for the model with the ionosphere have generally higher $\textnormal{Coh}^{2}$ values, especially for higher SH degress where the contribution of the ionosphere is expected to be larger. For instance, $C_{3}$ estimated from the $P_{3}^{2}$ mode boasts a squared coherence of $\mathrm{Coh}^{2}=0.95$ at the dominant period of $12$hrs, a drastic improvement over the mere $\mathrm{Coh}^{2}=0.42$ for the model without the ionosphere; $C_{4}$ estimated from the $P_{4}^{3}$ mode has $\mathrm{Coh}^{2}=0.78$ at the period of $8$hrs, compared to $\mathrm{Coh}^{2}=0.43$ for the model that neglects ionosphere (Fig. 19). Similar advantages can also be observed for $C_{2}$ estimated from the $P_{2}^{1}$ mode around the diurnal period band, and across almost the entire period range for $C_{3}$ and $C_{4}$, albeit both have lower coherences at longer periods. For the $C_{1}$ estimated from the first zonal SH $P_{1}^{0}$, we note that the transfer functions from both models deteriorate as they approach this band compared to longer periods, where the two estimates are nearly identical in terms of coherence.
@@
The difference in the estimated transfer functions and their coherence shows that a lack of proper treatment of the ionosphere in the model results in transfer functions that are wildly incoherent and physically inconsistent. Fortunately, this does not affect the $C_{1}$-response computed from the first zonal harmonic at periods longer than one day, rendering previous global reference models (e.g. Grayver et al., 2017; Kuvshinov et al., 2021) still valid. The situation is different for the 3-D mantle inversions. Previous studies (Velímský & Knopp, 2021; Kuvshinov et al., 2021) used external and internal coefficients derived from satellite and ground observations whereby the effect of the ionosphere was not explicitly modelled, likely resulting in many source effects being propagated to 3-D conductivity models. Therefore, parametrising the ionosphere in the model offers a path towards coherent and physical transfer functions, which are essential for probing the 3-D subsurface conductivity of the Earth with satellite and ground data.
Besides new opportunities for induction studies, our model also enables a much more flexible and versatile characterisation of the external magnetic field sources. Thanks to the relaxed assumptions on temporal behaviour of the fields and the environmental setting, our model is capable of exploiting data across all local times and during all magnetic conditions. To get a glimpse of the spacial structure and the temporal behaviour of the estimated external field, we visualise the results in terms of the streamfunction of an equivalent current, defined as
@@
During geomagnetic quiet times, our model recovers the dominant Sq current system in the ionosphere, as is shown in Fig. 20. The equivalent current of the Sq current system features one large-scale counter-clockwise vortex in the northern hemisphere, and one large-scale clockwise vortex in the southern hemisphere. Near the summer solstice, the foci of the Sq currents in the northern hemisphere are observed to consistently lag behind in local time compared to the southern foci, consistent with the previous observations on the seasonal variation in the local times of the focus positions (Yamazaki & Maute, 2017). Near the spring equinox, the northern and southern foci of the Sq currents alternate to lead in local times on a semi-diurnal basis.
@@
During geomagnetic storms, the estimated ionospheric current deviates considerably from a quiet Sq current, exemplified by the ionospheric equivalent current reconstructed during the geomagnetic storm on Sept. 8, 2017 (Fig. 21). Compared to the Sq current during quiet times, the equivalent sheet shows two distinct features during geomagnetic storms. First, the day-night dichotomy, ubiquitous during geomagnetic quiet times, is no longer present. Instead, the sheet current shows strong current amplitude across all local times (see UT=1:30 and 13:30 on 8 September). Second, the structurally robust vortices centred at mid-latitudes in the northern and southern hemispheres give way to multiple transient, rapid changing vortices centred at high latitudes (see UT=22:30, 1:30, 13:30, 16:30 on 8 September, when $\mathrm{Kp}\geq 7$). This may imply the energy input from the magnetosphere in high latitudes, which then propagates to lower latitudes by virtue of electric field penetration (Yamazaki & Maute, 2017). Interestingly, we see that in between the time bins with $Kp\geq 7$, the Sq current system still seems to persist, observable as the two vortices in the northern and southern hemispheres.
@@
Figure 21: Streamfunction of the equivalent current for the ionospheric field ($\Psi^{\mathrm{ion}}$) during a geomagnetic storm on Sept. 8, 2017. The same cutoff latitude of $70$ degrees is used. Note that the maximum value of the color scale is extended to twice of that for the quiet day plot (Fig. 20) to accommodate higher amplitudes of the streamfunction during the storm. Nevertheless, the strong equivalent currents saturate the colour scales when $\mathrm{Kp}>7$.
@@
Our model provides continuous description of the mid-latitude ionospheric and magnetospheric fields, covering all local times and magnetic conditions. The temporal resolution of the present model is three hours, indicating that shorter time-scale dynamics is not captured by the current model. The (in-)ability to capture these fast variations is controlled by a trade-off between space and time coverage that is dictated by the present distribution of observations. The new low-orbit geomagnetic satellites, such as the Macau Science Satellites (MSS) and a planned ESA NanoMagSat mission, will allow us to increase spatio-temporal resolution if the model. The use of spherical harmonics is another limiting factor in capturing localised current systems in polar and equatorial regions. Dedicated parametrisation of the polar electrojets based on loop or point currents can be integrated into our model to improve the fit at polar regions. Thanks to its flexibility, the model can be updated on the fly in real time, thus making it particularly suitable for working with low-time-lag datasets, such as the Swarm FAST data, in support of space weather operations and hazard assessment.

### v2

@@ 5.3 Fourier spectrum of external and induced fields
Our 10-year time series of the SH coefficients has a theoretical bandwidth of over four decades, i.e. $f_{\mathrm{Nyq}}/\Delta f>10^{4}$, where $f_{\mathrm{Nyq}}$ is the Nyquist frequency and $\Delta f$ is the frequency resolution. As such, it covers the entire period range from sub-diurnal to inter-annual periods. In this section, we explore the frequency content of the estimated fields by deriving their amplitude spectra. The spectra are computed directly as discrete Fourier transforms of the time series of the Gauss coefficients, using a forward normalisation that preserves the time-domain amplitude of a harmonic signal in the amplitude spectrum (i.e. a sinusoidal signal with 1 nT amplitude will result in a 1 nT peak at the corresponding frequency in the amplitude spectrum).
@@
The first zonal harmonic of the magnetospheric magnetic field, largely attributed to the ring current, is the most energetic contribution to the magnetic field of the external origin at periods longer than one day. The power of the spectrum for the first zonal harmonic quickly rolls off at periods shorter than one day, diurnal frequency bands, while exhibiting strong amplitude at longer periods (Fig. 17, upper panel). At longest periods, a prominent spectral peak can be observed between $2.0$ and $2.5$ years. This biennial oscillation has been previously observed in the solar magnetic field, and has been attributed to a harmonic of the solar cycle (Knaack & Stenflo, 2005). This notable feature can be similarly observed in the ionospheric signal (Fig. 17 lower panel) and in other degree-one modes (Fig. LABEL:fig:spec-q11-mi). The second, rather sharp peak in the magnetospheric field spectrum is at the semi-annual period (2 cycles per year). Since our estimation is conducted in the Earth-fixed frame, the semi-annual peak is likely associated with the axial tilt of the Earth and the associated seasonal variability in the Earth’s frame (Maus & Lühr, 2005; Olsen et al., 2005; Balasis & Egbert, 2006). A secondary, wider peak can be observed at the annual period (1 cycle per year). This can be a signature of the tail current in the magnetosphere (Maus & Lühr, 2005). As the $q_{10}^{\mathrm{mag}}$ component experiences semi-annual changes in the solar radiation, it is further modulated by the change in the Interplanetary Magnetic Field. Both annual and semi-annual peaks are observed not only in the first zonal harmonic component, but also in other dipole components of the magnetospheric field (Fig. LABEL:fig:spec-q11-mi upper panel). This can result from the transformation of the quasi-static magnetospheric field in GSM/SM frames to the Earth-fixed frame. In contrast, these peaks are weaker or not visible in the ionospheric spectrum (Figs. 17 and LABEL:fig:spec-q11-mi, lower panels), suggesting that ionospheric dynamics at $n=1$ is not as strongly modulated by the orbit-related periodicities. Note that this observation only holds for the $n=1$ spherical harmonic signals. The annual, semi-annual even quarterly peaks are clearly visible in both magnetospheric and ionospheric Gauss coefficients of $n>1$ (Figs. LABEL:fig:spec-s21-mi,LABEL:fig:spec-q32-mi).
@@
Figure 17: Amplitude spectra of the magnetospheric coefficient $q_{10}^{\mathrm{mag}}$ (upper panel), the ionospheric coefficient $q_{10}^{\mathrm{ion}}$ coefficient (middle panel) and the internally induced coefficient $g_{10}^{\mathrm{int}}$ (lower panel). Spectral peaks observed in the spectra are marked with blue triangles, with their peak periods annotated. Selected theoretical periods of the system are plotted as gray vertical lines for reference.
@@
In contrast to variations at long periods, the variations in diurnal period band are dominated by the ionospheric sources. One of the major features in the ionospheric signal in this period band is the diurnal mode and its harmonics. These are clearly visible as discrete spectral peaks at one, two, three and four cycles per day (cpd) (Fig. 18, lower panel). For the $P_{2}^{1}$ spatial mode, the fundamental time harmonic with a period of one day is the strongest, with an amplitude two orders of magnitudes above the average spectrum level. This mode, which has an azimuthal wavenumber $m=1$, and has a 1-cpd periodicity observed in the Earth-fixed geographic frame, represents a stationary $n=2$ mode in a solar-fixed frame (local-time term). This mode is predominantly influenced by the day-side structure of the Sq current system (Malin, 1973; Winch, 1981; Schmucker, 1999). The same holds also for the SH coefficients $s_{32}^{\mathrm{ion}}$ and $s_{43}^{\mathrm{ion}}$ with the azimuthal wavenumbers of $m=2$ and $m=3$, respectively. They exhibit strongest spectral peaks at 2 cpd (Fig. LABEL:fig:spec-s32-mi lower panel) and 3 cpd (Fig. LABEL:fig:spec-s43-mi lower panel), respectively. These correspond to the $n=3$ and $n=4$ modes that are stationary in a solar-fixed frame. The diurnal periods are also visible in the spectra of magnetospheric coefficients (Figs. 18, LABEL:fig:spec-s32-mi and LABEL:fig:spec-s43-mi upper panels), although they have a lower signal-to-noise ratio compared to the ionosphere.
Figure 18: Amplitude spectra of the magnetospheric coefficient $q_{21}^{\mathrm{mag}}$ (upper panel), the ionospheric coefficient $q_{21}^{\mathrm{ion}}$ coefficient (middle panel) and the internally induced coefficient $g_{21}^{\mathrm{int}}$ (lower panel) around the diurnal band. Spectral peaks observed in the spectra are marked with blue triangles, with their peak periods annotated. Selected theoretical periods of the system, including the diurnal period and its harmonics, and the periods for lunar daily variations $L_{p}$ with $p=1,2,3,4$ are plotted as gray vertical lines for reference.
@@
The estimated internal field contains signals that originate from both the magnetospheric and ionospheric field variations. For the $P_{1}^{0}$ component, for instance, the internal coefficient $g_{10}^{\mathrm{int}}$ inherits the strong semi-annual and Rieger-type spectral peaks as well as the peak at $29.7$ days and $25.5$ days from the magnetospheric signal, whereas these peaks are absent in the ionospheric signal (Fig. 17). For $n>1$ in diurnal bands, the induced coefficients show clear spectral peaks at $L_{p}$ tidal periods, which are only observed in the ionosphere but not in the reconstructed magnetospheric signal (Fig. 18). Additionally, our induced coefficients also contain oceanic tidal magnetic signals (Grayver et al., 2024). However, since we use both day-side and dark-side data and adopt a low-degree spherical harmonic representation of the field, the ocean tidal contributions are superposed by signals resulting from the ionospheric lunar daily variations (Malin & Chapman, 1970).
@@ 6 Discussion and Conclusions
Compared to the previous geomagnetic field models, our new modelling approach with the ionosphere parametrisation has several distinct features. First, the magnetic fields with ionospheric and magnetospheric origins are separated. This sets our model apart from the modelling approaches where the ionospheric field is neglected (Lesur et al., 2010; Finlay et al., 2020) or the magnetospheric field is removed by baseline removal and ring current corrections (Yamazaki & Maute, 2017; Guzavina et al., 2019). Second, no temporal behaviour is prescribed on either of the external fields. This is in contrast to many other models. For instance, the temporal behaviour for magnetospheric fields beyond $P_{1}^{0}$ is prescribed to be standing modes in SM or GSM coordinates in the CHAOS model (Finlay et al., 2020), or the ionospheric field is assumed to consist of daily and seasonal time-harmonics in the Comprehensive Model (Sabaka et al., 2018). Furthermore, the secondary induced field is co-estimated alongside the external fields, rather than modelled using an existing conductivity model, as is the case in Chulliat et al. (2016).
@@
When the geomagnetic field model fails to describe the ionospheric sources properly, the remnant ionospheric signal will be compensated by biasing the induced field and the magnetospheric field estimates. To test whether the induced and the magnetospheric fields are indeed biased if the ionospheric field is neglected, we need a physics-based metric to examine how the model with an ionosphere compares with a model without the ionosphere. For this purpose, we estimated the transfer functions (TFs) relating the induced and the external parts of the geomagnetic potential. In the frequency domain, assuming a radially symmetric electrical conductivity of the Earth, the Gauss coefficients of the induced field are related to those of the external field as
@@
where $\omega$ is the angular frequency, and $Q_{n}$ is the so-called $Q$-response (Schmucker, 1985), a functional of the subsurface conductivity $\sigma$. For a general 3-D subsurface conductivity, the relation above is not complete since a single external mode would induce many internal modes, all of which are coupled through a so-called Q-matrix (Olsen, 1999a). As the subsurface conductivity can be approximated to the first order as a radial conductivity profile, the diagonal elements of this matrix dominate, and the relation between the induced and external coefficients can to first order be approximated by the expression above. Therefore, the equation above is still a valid test since any realistic 3-D conductivity effects will be dwarfed by much stronger source effects.
@@
$\varepsilon_{nm}=\left\{\begin{aligned} &q_{n0},&&m=0\\ &\frac{1}{2}\left(q_{n|m|}-is_{n|m|}\right),&&m>0,\\ &\frac{1}{2}\left(q_{n|m|}+is_{n|m|}\right),&&m<0,\end{aligned}\right.\qquad\iota_{nm}=\left\{\begin{aligned} &g_{n0},&&m=0\\ &\frac{1}{2}\left(g_{n|m|}-ih_{n|m|}\right),&&m>0,\\ &\frac{1}{2}\left(g_{n|m|}+ih_{n|m|}\right),&&m<0.\end{aligned}\right.$ (20)
@@
The real part of the $C_{n}$-response is a depth of the maximum induced current, thus approximating the sounding depth at a given frequency (Weidelt, 1972). The value of $C_{n}$ at $\omega$ is estimated following the general method outlined in Olsen (1998). First, the time series of the complex Gauss coefficients $\varepsilon_{nm}$ and $\iota_{nm}$ (obtained from real counterparts using Eq. 20) are split into segments with length $3T$, where $T=2\pi/\omega$ is the period. Second, the short-time Fourier transform of the time series with Hamming time window is computed for every segment $\tau_{j}$ ($j=1,2,\cdots N_{\tau}$), generating the spectrum points within the segment, denoted as $\iota_{nm}(\tau_{j},\omega)$ and $\varepsilon_{nm}(\tau_{j},\omega)$. The final transfer function is estimated by performing a robust regression on the linear system (19), with $\varepsilon_{nm}(\tau_{j},\omega)$ and $\iota_{nm}(\tau_{j},\omega)$ estimated in the previous step. The resulting $C_{n}$-responses from the field estimates of the model with and without the ionosphere are visualised in Fig. 19 at $9$ log-spaced periods between $1.5$ days and $10$ days, as well as at the diurnal ($24$ hours), semidiurnal ($12$ hours) and terdiurnal ($8$ hours) periods. It is also possible to estimate these transfer functions at longer periods, but the effect of the ionosphere becomes smaller at periods longer than 20-30 days. To evaluate the linear correlation between the internal and external fields, we also computed the squared coherence ($\mathrm{Coh}^{2}$) for each spherical harmonic mode and each frequency, defined as
$\mathrm{Coh}^{2}=\frac{\left|\sum_{j=1}^{N_{\tau}}\iota^{*}_{nm}(\tau_{j},\omega)\,\varepsilon_{nm}(\tau_{j},\omega)\right|^{2}}{\sum_{j=1}^{N_{\tau}}\left|\iota_{nm}(\tau_{j},\omega)\right|^{2}\sum_{j=1}^{N_{\tau}}\left|\varepsilon_{nm}(\tau_{j},\omega)\right|^{2}}.$ (22)
@@
The TFs for the model with the ionosphere have generally higher squared coherence, especially for higher SH degress where the contribution of the ionosphere is expected to be larger. For instance, $C_{3}$ estimated from the $P_{3}^{2}$ mode boasts a squared coherence of $\mathrm{Coh}^{2}=0.95$ at the dominant period of $12$hrs, a drastic improvement over the mere $\mathrm{Coh}^{2}=0.42$ for the model without the ionosphere; $C_{4}$ estimated from the $P_{4}^{3}$ mode has $\mathrm{Coh}^{2}=0.78$ at the period of $8$hrs, compared to $\mathrm{Coh}^{2}=0.43$ for the model that neglects ionosphere (Fig. 19). Similar but somewhat smaller advantages can also be observed for $C_{2}$ estimated from the $P_{2}^{1}$ mode around the diurnal period band, and across almost the entire period range for $C_{3}$ and $C_{4}$, albeit both have lower coherences at longer periods. For the $C_{1}$ estimated from the first zonal SH $P_{1}^{0}$, we note that the transfer functions from both models deteriorate as they approach the diurnal band compared to longer periods, where the two estimates are nearly identical in terms of coherence.
@@
The difference in the estimated transfer functions and their coherence shows that a lack of proper treatment of the ionosphere in the model results in transfer functions that are in some cases significantly incoherent and physically inconsistent. Fortunately, this does not affect the $C_{1}$-response computed from the first zonal harmonic at periods longer than one day, rendering previous global reference models (e.g. Grayver et al., 2017; Kuvshinov et al., 2021) still valid. The situation is different for the 3-D mantle inversions. Previous studies (Velímský & Knopp, 2021; Kuvshinov et al., 2021) used external and internal coefficients derived from satellite and ground observations whereby the ionospheric effect was not explicitly modelled. The authors subtracted a model of the ionosphere (together with the corresponding induced field) with a prescribed periodic behaviour, constructed using quiet-time magnetic data as a part of the CI model (Sabaka et al., 2018). Failing to co-model ionospheric sources or extrapolating quiet-time ionosphere models to the entire data set will likely result in source effects being propagated to 3-D conductivity models. To illustrate this, we also first subtracted a CI-derived ionospheric model (with its induced field) from our dataset and then carried out the internal-external field separation. The resulting magnetospheric spectra show lunar tidal peaks (Fig. LABEL:fig:spec-ms4-MIOres), implying the ionospheric signal may have leaked into the magnetospheric estimates. At the same time, there is almost no improvement and even occasional deterioration in the coherence of the estimated TFs (Fig. LABEL:fig:Cn-10d-MIOres) compared to the internal-external field separation without the ionosphere on the original dataset (Fig. 19). Therefore, explicitly co-estimating the ionosphere is necessary to achieve coherent and physical transfer functions for probing the 3-D subsurface conductivity of the Earth with satellite and ground data.
Besides new opportunities for induction studies, our model also enables a more flexible and versatile characterisation of the external magnetic field sources. Thanks to the relaxed assumptions on temporal behaviour of the fields and the environmental setting, our model is capable of exploiting data across all local times and during all magnetic conditions. To get a glimpse of the spacial structure and the temporal behaviour of the estimated external field, we visualise the results in terms of the streamfunction of an equivalent current, defined as
@@
During geomagnetic quiet times, our model recovers the dominant Sq current system in the ionosphere, as is shown in Fig. 20. The equivalent current of the Sq current system features one large-scale counter-clockwise vortex in the northern hemisphere, and one large-scale clockwise vortex in the southern hemisphere. Near the summer solstice, the foci of the Sq currents in the northern hemisphere are observed to consistently lag behind in local time compared to the southern foci, consistent with the previous observations on the seasonal variation in the local times of the focus positions (Yamazaki & Maute, 2017). Near the spring equinox, the northern and southern foci of the Sq currents alternate to lead in local times on a semi-diurnal basis. These seasonal changes of the Sq system are also reflected by the annual, semi-annual and seasonal periodicities in the ionospheric Gauss coefficient spectra at spherical harmonic degree $n>1$ (Figs. LABEL:fig:spec-s21-mi and LABEL:fig:spec-q32-mi lower panels).
@@
During geomagnetic storms, the estimated ionospheric current deviates considerably from a quiet Sq current, exemplified by the ionospheric equivalent current reconstructed during the geomagnetic storm on Sept. 8, 2017 (Fig. 21). Compared to the Sq current during quiet times, the equivalent sheet shows two distinct features during geomagnetic storms. First, the day-night dichotomy, ubiquitous during geomagnetic quiet times, is no longer present. Instead, the sheet current shows strong current amplitude across all local times (see UT=1:30 and 13:30 on 8 September). Second, the structurally robust vortices centred at mid-latitudes in the northern and southern hemispheres give way to multiple transient, rapid changing vortices centred at high latitudes (see UT=22:30, 1:30, 13:30, 16:30 on 8 September, when $Kp\geq 7$). This may imply the energy input from the magnetosphere in high latitudes, which then propagates to lower latitudes by virtue of electric field penetration (Yamazaki & Maute, 2017). Interestingly, we see that in between the time bins with $Kp\geq 7$, the Sq current system still seems to persist, observable as the two vortices in the northern and southern hemispheres.
@@
Figure 21: Streamfunction of the equivalent current for the ionospheric field ($\Psi^{\mathrm{ion}}$) during a geomagnetic storm on Sept. 8, 2017. The same cutoff latitude of $70$ degrees is used. Note that the maximum value of the color scale is extended to twice of that for the quiet day plot (Fig. 20) to accommodate higher amplitudes of the streamfunction during the storm. Nevertheless, the strong equivalent currents saturate the colour scales when $Kp>7$.
Our model provides a continuous description of the mid-latitude ionospheric and magnetospheric fields for the period 2014-2023, covering all local times and magnetic conditions. The temporal resolution of the presented model is three hours, indicating that dynamics on a time scale that is similar to or less than three hours cannot be well captured. This is especially pronounced in phases of geomagnetic storms with fast Dst variations, where we observe higher misfits of the model ($R^{2}\sim 0.7$ for UT=22:30, 01:30, 13:30, Fig. 21). In comparison, the misfits during other times tend to be lower ($R^{2}>0.85$ for all other snapshots). In addition, the spatial parametrisation of the current model is up to SH degree and order $4$ for induced and magnetospheric fields, and up to degree and order $5$ for ionospheric field. Localised features, such as equatorial electrojets and the polar current system, cannot be described within our parametrisation and hence have been avoided in data selection. The (in-)ability to capture certain temporal and spatial variations of the magnetic fields is controlled by a trade-off between space and time coverage of the data, which is ultimately dictated by the distribution of the observations.
@@
In light of these limitations, our framework can benefit both from the incorporation of additional data, as well as dedicated parametrisation for localised current systems. The new low-orbit geomagnetic satellites, such as the Macau Science Satellites (MSS) and a planned ESA NanoMagSat mission, will significantly improve the spatial (especially local-time) coverage, allowing us to increase the spatio-temporal resolution of the model. Dedicated parametrisation of polar electrojets based on loop or point currents can be integrated into our model to improve the fit at polar regions. Thanks to its flexibility, the model can be updated on the fly in real time, thus making it particularly suitable for working with low-time-lag datasets, such as the Swarm FAST data, in support of space weather operations and hazard assessment.